### v1

# The spectral genus of an isolated hypersurface singularity and a conjecture relating to the Milnor number

Abstract.
In this paper, we introduce the notion of spectral genus $\widetilde{p}_{g}$ of a deformation of a germ of an isolated hypersurface singularity $(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, defined as a sum of small exponents of monodromy eigenvalues. The number of these is equal to the geometric genus $p_{g}$, and hence $\widetilde{p}_{g}$ can be considered as a secondary invariant to it. We then explore a secondary version of the Durfee conjecture on $p_{g}$, regarding the relationship between $\widetilde{p}_{g}$ the Milnor number $\mu$, to the effect that
$\frac{\mu}{(n+2)!}>\widetilde{p}_{g}.$
We provide evidence by confirming the conjecture in several cases, including homogeneous singularities and singularities with large Newton polyhedra, and quasi-homogeneous or irreducible curve singularities.
Our conjecture is shown to relate closely to the asymptotic behavior of the holomorphic analytic torsion of the sheaf of holomorphic functions on a degeneration of projective varieties, potentially indicating deeper geometric and analytic connections.
Key words and phrases:
Isolated hypersurface singularities, Spectrum of singularities, Milnor numbers
2010 Mathematics Subject Classification:
Primary: 32S25, 32S30. Secondary: 14M25, 32S20, 58J52.
1. Introduction
1.1.
In 1978, Durfee conjectured an inequality between the geometric genus $p_{g}$ of complete intersection isolated surface singularities and their Milnor numbers $\mu$, in [11]. While the initial expectation does not hold in such generality, the conjecture was later generalized in [27] by K. Saito to a conjecture about general germs of $n$-dimensional isolated hypersurface singularities [^1] In this paper, by a singularity, we mean a non-regular point., defined by germs of holomorphic functions $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, as an inequality
(1.1) $\frac{\mu}{(n+1)!}>p_{g}.$
This generalized inequality appeared naturally in his investigation of the distribution of the spectrum of the semi-simple part of the monodromy in the same article. Particular cases and variants of the Durfee-type conjecture (1.1) have been established by Yau–Zhang [36] and Kerner–Némethi [17, 18].
1.2.
By work of M. Saito and Steenbrink [28, 33], the geometric genus of an isolated hypersurface singularity $f=0$ is related to the cohomology of the Milnor fiber and its mixed Hodge structure by $p_{g}=\dim\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})$. We define the spectral genus of the singularity (cf. §2.2 below) as
$\widetilde{p}_{g}=\sum\lambda_{j},$
where the sum is over the rational numbers $\lambda_{j}\in[0,1)$ such that $\exp(2\pi i\lambda_{j})$ is an eigenvalue of the semi-simple part of the monodromy acting on $\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})$. It is a sort of secondary invariant of the geometric genus, and is in particular zero for rational singularities. One may wonder if there is a corresponding secondary version of Durfee’s conjecture. In light of this, we propose the following:
Conjecture.
For an $n$-dimensional isolated hypersurface singularity, we have an inequality:
  1. (1)
    (Weak form)
    $\frac{\mu}{(n+2)!}>\widetilde{p}_{g}.$
  2. (2)
    (Strong form)
    $\frac{\mu}{(n+2)!}\geq\widetilde{p}_{g}+\frac{1}{(n+2)!}.$
The inequality in the strong form is attained for ordinary double points, since in this case the spectral genus is zero and the Milnor number is one. It seems natural to wonder if this is the only case where the inequality is attained. Moreover, if K. Saito’s conjectures in [27] about the distribution of monodromy eigenvalues hold, then the above inequalities are essentially optimal. Namely, it implies (cf. Proposition 7.2) that
$\frac{\widetilde{p}_{g}}{\mu}\to\frac{1}{(n+2)!}$
as the singularities get worse.
1.3.
Our contribution is the following partial confirmation:
Main Theorem.
The strong form of the conjecture is true in the following cases:
  1. (i)
    Homogeneous singularities in arbitrary dimension.
  2. (ii)
    Quasi-homogeneous curve singularities.
  3. (iii)
    Irreducible curve singularities.
  4. (iv)
    Generic and convenient singularities with large Newton polyhedra in arbitrary dimension.
For the proofs, see the corresponding sections in the article, relying in one way or another on the fact that the spectrum for non-degenerate singularities can be described in terms of Newton polyhedra. The case of irreducible curve singularities covers also degenerate cases.
We remark that other singularities have been tested with the software Singular [9], allowing instantaneous computation of all involved invariants. In fact, the conjecture was initially based on an extensive numerical investigation, utilizing this software, of the expressions $\mu/6-\widetilde{p}_{g}$ for curve singularities and $\mu/24-\widetilde{p}_{g}$ for surface singularities.
1.4.
The extension of the conjecture to higher dimensions was partially motivated by the asymptotic behavior of the holomorphic analytic torsion, which is defined in terms of regularized determinants of Dolbeault–Laplace operators. More precisely, the behavior of $\tau(\mathcal{O}_{\mathcal{X}_{t}})$ for a degenerating family of projective varieties $\mathcal{X}\to\Delta$ with isolated singularities, is captured by the expression $(-1)^{n}(\mu/(n+2)!-\widetilde{p}_{g})$ (cf. Corollary 3.2). This is related to the work of Yoshikawa on the singularities of the Quillen metric [37, 38]. See also [12].
The weak form of the conjecture is connected to the question whether the function
(1.2) $X\mapsto\tau(\mathcal{O}_{X})^{(-1)^{n}}$
is continuous over the space of hypersurfaces in $\mathbb{P}^{n+1}$. More precisely, the weak form of the conjecture indicates that the analytic torsion even vanishes along the discriminant locus. In fact, this continuity statement implies the weak form of the conjecture (cf. §2.7 and §3.7). This is based on the fact the function in (1.2), if continuous, vanishes on the locus of ordinary double points because the weak form of the conjecture is known then, and any isolated singularity is a limit of ordinary double points. This formulation of the problem possibly invites analytic techniques, of global nature, to study the conjecture, which is of local type.
1.5.
In future work we aim to apply and extend part of this discussion to the asymptotic behavior of the BCOV invariant of Calabi–Yau varieties, introduced in dimension 3 in [15] and in general dimension in [13]. Some instances of such asymptotics were used to establish new cases of genus one mirror symmetry in [14]. Further qualitative discussions of the same type could provide even further cases and new insights into mirror symmetry phenomena.
2. Invariants of isolated singularties
In this section, to set up notation and for the convenience of the reader, we recall some classical invariants of isolated hypersurface singularities. We will be considering the germ of a holomorphic function $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ defining an isolated hypersurface singularity at the origin, and we will suppose that $n\geq 1$. Somewhat abusively, in this text we will simply write $f=0$.
2.1.
It is convenient to study these isolated hypersurface singularities through the cohomology of the Milnor fiber $H^{n}(\operatorname{Mil}_{f})$. Note that the (reduced) cohomology of the Milnor fiber vanishes in other degrees. Its dimension over $\mathbb{C}$ is the Milnor number of the singularity:
$\mu=\dim\mathbb{C}\{z_{0},\ldots,z_{n}\}/\left(\partial f/\partial z_{0},\ldots,\partial f/\partial z_{n}\right).$
The cohomology group $H^{n}(\operatorname{Mil}_{f})$ is equipped with a canonical mixed Hodge structure by [32]. We denote the corresponding Hodge filtration by $F$. Moreover, the semi-simple part of the monodromy, denoted by $T_{s}$, acts on this mixed Hodge, in the sense that it preserves the weight and Hodge filtrations [32, Theorem 4.1].
2.2.
Consider the the genus of the singularity $V=\{f=0\}$:
(2.1) $p_{g}=\dim\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})=\begin{cases}\dim R^{n-1}\pi_{\ast}\mathcal{O}_{X}&\mbox{if }n>1,\\ \dim\pi_{\ast}\mathcal{O}_{X}/\mathcal{O}_{V}&\mbox{if }n=1,\end{cases}$
where $\pi:X\to V$ is a desingularization. The equality of the various quantities is proven in [28, Theorem 1] and [33, Proposition 2.13]. Recall from the introduction that we likewise define the spectral genus of the singularity, as the expression
(2.2) $\widetilde{p}_{g}=\sum\lambda_{j},$
where the sum is over all $10<\lambda_{j}<$ 1 such that $\exp(2\pi i\lambda_{j})$ is an eigenvalue of $T_{s}$ acting on $\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})$. More compactly, (2.2) can be recast as
(2.3) $\widetilde{p}_{g}=\frac{1}{2\pi i}\operatorname{tr}\left(\log(T_{s})\mid{\operatorname{Gr}^{n}_{F}H^{n}(\operatorname{Mil}_{f})}\right),$
where $\log$ is the branch of the logarithm whose imaginary part lies in $[0,2\pi)$.
2.3.
We describe the relationship between the spectral genus and the spectrum of an isolated hypersurface singularity. We first follow Steenbrink’s presentation in [34, Section 2] and [35, Section 1], and we adopt his conventions. See also [25, Section 12.1.3]. The spectral numbers are associated to the triple $(H^{n}(\operatorname{Mil}_{f}),F^{\bullet},T_{s})$. These are rational numbers $\alpha$, given with multiplicities, uniquely determined by the following conditions:
  1. (1)
    $\exp(-2\pi i\alpha)$ is an eigenvalue of $T_{s}$ acting on $\operatorname{Gr}^{p}_{F}H^{n}(\operatorname{Mil}_{f})$, for some $p=0,\ldots,n$.
  2. (2)
    For $\alpha$ and $p$ as in the first point, $p=[n-\alpha]$. Equivalently, we take $n-p-1<\alpha\leq n-p$.
  3. (3)
    The multiplicity of $\alpha$ is the multiplicity of $\exp(-2\pi i\alpha)$.
Denote by $\{\alpha_{j}\}_{j=1,\ldots,\mu}$ the collection of spectral numbers, with multiplicities. This collection is invariant under $\alpha\mapsto n-1-\alpha$. The spectral numbers hence belong to the interval $(-1,n)$. With this understood, we see that the spectral genus $\widetilde{p}_{g}$ can be expressed as
$\widetilde{p}_{g}=-\sum_{-1<\alpha_{j}<0}\alpha_{j}.$
Some authors shift the spectral numbers by one, so that they are given in the form $\alpha^{\prime}_{j}=\alpha_{j}+1$ and belong to $(0,n+1)$. With this convention, we have
(2.4) $\widetilde{p}_{g}=\sum_{0<\alpha_{j}^{\prime}<1}(1-\alpha_{j}^{\prime}).$
For instance, this is the case of the works of M. Saito [29, 30] used below. Depending on the context, one convention may be more adapted than the other, and we will use both.
For later use, we recall the definition of the spectral polynomial associated to $f$, which for the spectral numbers taken in $(0,n+1)$ is given by
(2.5) $\mathrm{Sp}_{f}(T)=\sum_{j}T^{\alpha_{j}^{\prime}}\in\mathbb{Z}[T^{\mathbb{Q}}].$
2.4.
The Milnor number $\mu$ and the geometric genus $p_{g}$ depend only on the fiber $V=f^{-1}(0)$, but not on the chosen deformation [34, Theorem 2.9]. The spectral genus depends on the whole germ $f$. Nevertheless, the spectrum is constant in any deformation of isolated hypersurface singularities with constant $\mu$ [34, Theorem 2.8], and hence so is the spectral genus.
2.5.
Consider a connected complex manifold $\mathcal{X}$ of dimension $n+1$, and a flat, projective morphism $g:\mathcal{X}\to\Delta$, which is a holomorphic submersion outside the origin. We will suppose that the central fiber has at most isolated singularities. If $x_{i}\in\mathcal{X}_{0}$ is such a singular point, the germ $(\mathcal{X},x_{i})\to(\Delta,0)$ is isomorphic to some $f_{i}:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$ and admits a Milnor fiber $\operatorname{Mil}_{f_{i}}$. In this setting, we denote by $\mu$ and $\widetilde{p}_{g}$ the sum of the Milnor numbers and spectral genera of the singularities $x_{i}$, respectively. Therefore, there is a decomposition
$\frac{\mu}{(n+2)!}-\widetilde{p}_{g}=\sum_{i}\left(\frac{\mu_{i}}{(n+2)!}-\widetilde{p}_{g,i}\right).$
2.6.
The cohomologies of the Milnor fibers of the singularities $x_{i}$ sit in an exact sequence
(2.6) $0\to H^{n}(\mathcal{X}_{0})\to H^{n}(\mathcal{X}_{t})\to\bigoplus_{i}H^{n}(\operatorname{Mil}_{f_{i}})\to H^{n+1}(\mathcal{X}_{0})\to H^{n+1}(\mathcal{X}_{t})\to 0,$
and for $q\neq n,n+1$ there is an isomorphism
(2.7) $H^{q}(\mathcal{X}_{0})\simeq H^{q}(\mathcal{X}_{t}),$
where $t\in\Delta\setminus\{0\}$. When $H^{n}(\mathcal{X}_{0})$ is given the canonical mixed Hodge structure of Deligne, and $H^{n}(\mathcal{X}_{t})$ is given the limit mixed Hodge structure, this is moreover an exact sequence of mixed Hodge structures. It is equivariant with respect to the semi-simple part of the monodromy, $T_{s}$. For details, we refer to Steenbrink [32, Section 3.3] and Navarro Aznar [23, Section 14].
2.7.
If we start with the germ of an isolated singularity $f:(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, it admits a good compactification $g:\mathcal{X}\to\Delta$. By this, we mean:
  1. (1)
    We are given a complex manifold $\mathcal{X}$, and a flat, projective morphism $g:\mathcal{X}\to\Delta$, which is a submersion outside the origin. We refer to $\mathcal{X}\to\Delta$ as a degeneration of projective varieties, or simply a projective degeneration.
  2. (2)
    The special fiber $\mathcal{X}_{0}$ has only one isolated singularity $x$.
  3. (3)
    There is an open subset $U$ of $x$ such that $(g|_{U},x)\to(\Delta,0)$ is isomorphic to $f$.
By an argument of Brieskorn [6, Section 1.1], such a compactification exists, and one can further suppose that it is given by a family of hypersurfaces in $\mathbb{P}^{n+1}$. This moreover shows that properties such as the positivity of $\frac{\mu}{(n+2)!}-\widetilde{p}_{g}$ can be studied equivalently for proper families or in the local setting.
3. Analytic torsion of $\mathcal{O}_{X}$
In this section, we discuss the asymptotic behavior of the holomorphic analytic torsion of the sheaf of holomorphic functions, for a degeneration of projective varieties.
3.1.
Let $X$ be a compact analytic space and suppose that we are given a holomorphic vector bundle $E$ on $X$. The determinant of the cohomology is the line
(3.1) $\lambda(E)=\bigotimes_{q}\det H^{q}(X,E)^{(-1)^{q}}.$
More generally, for a flat proper morphism of complex analytic spaces $\mathcal{X}\to S$ over a complex analytic manifold $S$, and a vector bundle $E$ on $\mathcal{X}$, there is a line bundle $\lambda(E)$ on $S$, whose fibers over $s\in S$ are given by $\lambda(E|_{\mathcal{X}_{s}})$, see [3, Section 4.1]. It is also referred to as the determinant of the cohomology. If $\mathcal{X}\to S$ is the analytification of an algebraic family, this construction is the analytification of the Knudsen–Mumford determinant [19]. It has a natural grading, which for the purposes of this article, together with various sign issues, we can ignore.
3.2.
Let $X$ be a compact Kähler manifold. The analytic torsion of a hermitian vector bundle $E$ on $X$ is given by
$\tau(X,E)=\exp\left(\sum(-1)^{q+1}q\zeta^{\prime}_{0,q}(0)\right)$
where, for $\operatorname{Re}(s)\gg 0$, $\zeta_{0,q}(s)$ is given by
$\zeta_{0,q}(s)=\sum\frac{1}{\mu_{j}^{s}},$
whose sum runs over positive eigenvalues $\mu_{j}$ of the Dolbeault-Laplacian $\Delta_{\overline{\partial}}$ acting on $A^{0,q}(E)$. It depends on both the Kähler metric on $X$ and the hermitian metric on $E$.
3.3.
Let $X$ be a Kähler manifold and $E$ a hermitian vector bundle on $X$. The determinant of the cohomology in (3.1) is equipped with two metrics, the $L^{2}$-metric and the Quillen metric. The $L^{2}$-metric $h_{L^{2}}$ on $\lambda(E)$ is defined by representing the Dolbeault cohomology groups $H^{q}(X,E)$ by harmonic forms, and using the natural metric from Hodge theory on $A^{0,q}(E)$-forms. The Quillen metric on $\lambda(E)$ is defined by
(3.2) $h_{Q}=h_{L^{2}}\cdot\tau(X,E),$
where $\tau(X,E)$ is the analytic torsion described in (3.2).
For a family of Kähler manifolds $\mathcal{X}\to S$ and a hermitian vector bundle $E$, the Quillen metric varies smoothly. If the dimensions $s\mapsto h^{q}(\mathcal{X}_{s},E|_{\mathcal{X}_{s}})$ are constant on $S$ for all $q$, the $L^{2}$-metric is also smooth.
3.4.
Let $\mathcal{X}\to\Delta$ be a degeneration of projective varieties of dimension $n$ as in §2.7 (1), with at most isolated singularities in the central fiber $\mathcal{X}_{0}$. Suppose that $\mathcal{X}$ is equipped with a hermitian metric, which is fiberwise Kähler, and that we are given a hermitian vector bundle $E$ on $\mathcal{X},$ of rank $e$. Let $\sigma$ be a holomorphic trivialization of the determinant of the cohomology $\lambda(E)$. By the main result of [37], we have
(3.3) $\log\|\sigma\|_{Q}^{2}=\frac{(-1)^{n}}{(n+2)!}\mu\cdot e\cdot\log|t|^{2}+o(\log|t|),\quad\text{as}\quad t\to 0.$
3.5.
Suppose now that $\mathcal{X}\to\Delta$ is as in §3.4, but $\mathcal{X}$ is actually equipped with a Kähler metric, whose associated Kähler form is rational when restricted to smooth fibers.
Proposition 3.1.
Let $\sigma$ be a trivialization of $\lambda(\mathcal{O}_{\mathcal{X}})$. Then, for small $t$,
$\log\|\sigma\|^{2}_{L^{2}}=(-1)^{n}\widetilde{p}_{g}\log|t|^{2}+o(\log(t)).$
Proof.
First of all, the determinant of the cohomology of the sheaf of holomorphic functions is invariant under blowups in the special fiber, because the total space is smooth and hence has only rational singularities. We can hence suppose the central fiber of $\mathcal{X}\to\Delta$ has normal crossings. In this case, by [13, Theorem C] we find that
$\log\|\sigma\|_{L^{2}}^{2}=\left(\sum(-1)^{q}\alpha^{0,q}\right)\log|t|^{2}+o(\log|t|),$
where $\alpha^{0,q}$ is defined as minus (the lower extension of) the logarithm of the semi-simple part of the monodromy acting on $H^{0,q}(\mathcal{X}_{t})$. One can infer directly from the definition that $\alpha^{0,n}=\widetilde{p}_{g}$. Moreover, $\alpha^{0,q}=0$ otherwise. This latter fact follows from (2.6) and (2.7), paired with the fact that $H^{q}(\mathcal{X}_{0})$ has trivial monodromy for any $q$. ∎
3.6.
Let $\mathcal{X}\to\Delta$ be a degeneration as in §3.5. We conclude by §3.3 and Proposition 3.1 the following:
Corollary 3.2.
The analytic torsion of the sheaf of holomorphic functions, endowed with the trivial metric, has the following asymptotic behavior for $t$ close to 00:
$\log\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})=(-1)^{n}\left(\frac{\mu}{(n+2)!}-\widetilde{p}_{g}\right)\log|t|^{2}+o(\log|t|).$
In particular, if the weak form of the conjecture holds, then $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}$ goes to zero as $t$ goes to 0.
Notice that, while the analytic torsion depends on the choice of metrics, the asymptotic behavior is a topological invariant. Also, for families of curves, the $L^{2}$-norm is independent of the choice of auxiliary metric, and the asymptotic behavior is then valid for any choice of Kähler metric on the total space $\mathcal{X}$.
3.7.
A situation where the corollary applies is that of a degeneration of projective varieties endowed with an embedding $\mathcal{X}\subseteq\mathbb{P}^{M}\times\Delta$ over $\Delta$, and such that $\mathcal{X}$ is equipped with a smooth Kähler metric $\omega$ on $\mathcal{X}$ whose cohomology class $[\omega_{t}]$ for $t\in\Delta\setminus\{0\}$ is given by the canonical polarization coming from $\mathbb{P}^{M}$.
A variant of the above arises from the tautological family $\mathcal{H}$ of hypersurfaces of degree $d$ in $\mathbb{P}^{n+1}$ over the space of parameters $\mathbb{P}^{N}$, where $N={n+1\choose d}-1$. Then $\mathcal{H}$ is a smooth space and inherits a Kähler metric from $\mathbb{P}^{n+1}$ and $\mathbb{P}^{N}$. Hence, the restriction along any curve $\Delta\to\mathbb{P}^{N}$ in which the total space is still smooth is an example of such a degeneration.
4. Quasi-homogeneous singularities
In this section, we address several cases of the conjecture in the setting of quasi-homogeneous singularities. We also recall some useful notions about Newton polyhedra also utilized in later sections.
4.1. Newton polyhedra
Let $f(x_{0},\ldots,x_{n})=\sum_{k\in\mathbb{N}^{n+1}}a_{k}x^{k}$ be a power series with complex coefficients, with $a_{0}=0$ and where we define $x^{k}=x_{0}^{k_{0}}\cdots x_{n}^{k_{n}}$. The Newton diagram of the singularity is constituted of the following polyhedra. The upper Newton polyhedron associated to $f$, denoted by $\Gamma_{+}$, is the convex hull of the set $\bigcup_{a_{k}\neq 0}(k+\mathbb{R}^{n+1}_{+})$. The associated Newton boundary, denoted by $\Gamma$, is the boundary of $\Gamma_{+}$. We denote by $\Gamma_{c}$ the compact Newton boundary, meaning the union of the compact faces of $\Gamma$. The lower Newton polyhedron, denoted by $\Gamma_{-}$, is the union of the lines joining the origin with the points on $\Gamma_{c}$. Since we only deal with lower Newton polyhedra, we will usually refer to these as simply Newton polyhedra.
Below we display a Newton diagram. In the picture, the lower Newton polyhedron is determined by the vertices $(0,0)$, $A$, $B$, $C$, $D$, and it is delimited by the segments painted in red. The compact Newton boundary has three faces, namely $AB$, $BC$, $CD$. The whole Newton boundary has two more unbounded faces, painted in blue. The upper Newton polyhedron is the region above the Newton boundary.
[…]
4.2.
Let $n\geq 1$ be an integer, and consider a quasi-homogeneous polynomial $f(x_{0},\ldots,x_{n})$, with rational weights $w_{0},\ldots,w_{n}>0$. This means that for any $\lambda\in\mathbb{C}$, we have $f(\lambda^{w_{0}}x_{0},\ldots,\lambda^{w_{n}}x_{n})=\lambda\cdot f(x_{0},\ldots,x_{n})$. We suppose
$f(x_{0},\ldots,x_{n})=0$
has an isolated singularity at the origin. Then, the weights satisfy $w_{i}<1$.
For quasi-homogeneous singularities, the Milnor number and the spectral genus depend only on the weights, and are given by the following formulas:
(4.1) $\mu=\prod_{i}\left(\frac{1}{w_{i}}-1\right)$
and
(4.2) $\widetilde{p}_{g}=\sum\left(1-k_{0}w_{0}-\cdots-k_{n}w_{n}\right),$
where the sum runs over integers $k_{i}>0$ such that $\sum k_{i}w_{i}<1$. These are the interior lattice points of the Newton polyhedron of the polynomial $f(x_{0},\ldots,x_{n})$. Eventually, for convenience, we may include the lattice points with $\sum k_{i}w_{i}=1$, since these contribute zero to the sum. For the Milnor number, the formula (4.1) is given in [21, Theorem 1]. The expression (4.2) for $\widetilde{p}_{g}$ can be derived from the computation of the spectral polynomial of $f$ (cf. (2.5) for the definition), due to Steenbrink [32, Example 5.11]:
$\mathrm{Sp}_{f}(T)=\prod_{j}\frac{T^{w_{j}}-T}{1-T^{w_{j}}}.$
To this end, it is enough to expand the polynomial into a power series of $T$ with rational exponents, and collect the terms whose exponents are strictly smaller than one. For this purpose, one can ignore the negative $T$ in the numerator, and expand $T^{w_{j}}/(1-T^{w_{j}})=T^{w_{j}}+T^{2w_{j}}+\ldots$. One finds that the exponents $\alpha$ strictly less than one are of the form $\alpha=\sum k_{i}w_{i}$, where the $k_{i}$ run over all the possible integers $k_{i}>0$ such that $\alpha<1$.
4.3. Homogeneous polynomials
It would be interesting to establish the conjecture for Brieskorn–Pham singularities of the form
$f(x)=x_{0}^{a_{0}}+x_{1}^{a_{1}}+\cdots+x_{n}^{a_{n}}=0,$
corresponding to the choices of weights $w_{i}=1/a_{i}$. To this end, we treat the case when $f$ be a homogenenous polynomial in arbitrary dimension, so that all $w_{i}=\frac{1}{d}$. In this case (4.1) gives
(4.3) $\mu=(d-1)^{n+1}.$
The spectral genus can be computed explicitly, based on the following elementary identity:
Lemma 4.1.
Let $d>1$ be an integer, and $n\geq 1$. Then
(4.4) $\sum\left(d-k_{0}-\cdots-k_{n}\right)=\frac{d(d-1)\cdots(d-(n+1))}{(n+2)!},$
where the sum runs over integers $k_{i}>0$, such that $\sum k_{i}<d$. Consequently,
$\widetilde{p}_{g}=\frac{(d-1)\cdots(d-(n+1))}{(n+2)!}.$
Proof.
An inductive argument shows that the sum is a polynomial of degree $n+2$ in $d$. Moreover, the sum is empty for $d=0,\ldots,n+1,$ so that the polynomial is of the form $C\cdot d(d-1)\cdots(d-(n+1))$ for some constant $C$. Since the sum is equal to $1$ for $d=n+2$ one sees that $C=\frac{1}{(n+2)!}$. ∎
The following proposition follows immediately from (4.3) and Lemma 4.1:
Proposition 4.2.
Let $f=0$ define an isolated degree $d$ homogeneous singularity at the origin. Then the strong form of the conjecture is true. Moreover, for a fixed $n$,
$\frac{\widetilde{p}_{g}(d)}{\mu(d)}\nearrow\frac{1}{(n+2)!},\quad\hbox{ as }\quad d\to+\infty,$
where $\mu(d)$ and $\widetilde{p}_{g}(d)$ denote the corresponding Milnor number and spectral genus.
4.4. Quasi-homogeneous singularities in dimension one
The purpose of the rest of this section is to prove the following theorem:
Theorem 4.3.
When $n=1$, the strong form of the conjecture is true in the case of quasi-homogeneous singularities.
In preparation for the proof, we first notice that the invariants $\mu$ and $\widetilde{p}_{g}$ only depend on the analytical type of the germ of the singularity. By [26, Satz 1.3] we can assume that $w_{i}\leq\frac{1}{2}$. Then, by [39], the weights depend only on the topological type of the singularity, which is one of the following:
(4.5) $x^{a}+y^{b}=0,$
(4.6) $x(x^{a}+y^{b})=0,$
(4.7) $xy(x^{a}+y^{b})=0.$
In the following sections, we will prove the theorem by analyzing these cases.
4.5.
We begin our treatment of the singularities in §4.4 by the following computation, which is due to Mordell [22] in the case when $a$ and $b$ are relatively prime:
Proposition 4.4.
Let $a,b\geq 2$ be two integers. Define $k=\gcd(a,b)$ and write $a=k\cdot a^{\prime},b=k\cdot b^{\prime}$. Then we have
(4.8) $\sum\left(1-\frac{x}{a}-\frac{y}{b}\right)=\frac{(a-1)(b-1)}{6}-\frac{a^{\prime}+b^{\prime}}{12}(k-1)-\frac{(a^{\prime}-1)(b^{\prime}-1)(a^{\prime}+b^{\prime}+1)}{12a^{\prime}b^{\prime}},$
where the sum runs over the interior lattice points of the triangle with vertices $(0,0)$, $(a,0)$, $(0,b)$.
Proof.
The case when $a,b$ are relatively prime follows immediately from the computation of Mordell [22, Equation (4)]. The sum in (4.8) is
(4.9) $\frac{(a-1)(b-1)}{6}-\frac{(a-1)(b-1)(a+b+1)}{12ab}.$
In general, we describe the main steps of the reasoning and leave the details to the reader. Let $k=\gcd(a,b)$, and decompose $a=a^{\prime}k$ and $b=b^{\prime}k$. We cut the triangle into smaller pieces, as in the following picture (for which $k=4$):
[…]
In the picture, there are $k$ small triangles obtained by appropriately translating the triangle of vertices $(0,0)$, $(a^{\prime},0)$, $(0,b^{\prime})$. The rest is divided into rectangles. After taking into account the appropriate translations, the evaluation of the sum in (4.8) restricted to the interior points of the smaller triangles reduces to Mordell’s computation (4.9). The evaluation of the sum (4.8) on the interior lattice points of the rectangles is elementary and reduces to some double sums of consecutive integers. Then we add the contribution from the dashed lines, which are dealt with in the same way. This concludes the proof. ∎
Proposition 4.5.
The strong form of the conjecture is true for the singularities of the form (4.5).
Proof.
Recall by (4.2) that $\widetilde{p}_{g}$ is given by the sum in (4.8). We first assume that $a$ and $b$ are relatively prime. Without loss of generality, we may suppose that $a\geq 2$ and $b\geq 3$. Then, by the formula of Proposition 4.4 :
$\frac{\mu}{6}-\widetilde{p}_{g}=\frac{(a-1)(b-1)(a+b+1)}{12ab}\geq\frac{1}{6}.$
In the general case, the same formula with $k\geq 2$, combined with the value of the Milnor number, gives
$\frac{\mu}{6}-\widetilde{p}_{g}=\frac{a^{\prime}+b^{\prime}}{12}(k-1)+\frac{(a^{\prime}-1)(b^{\prime}-1)(a^{\prime}+b^{\prime}+1)}{12a^{\prime}b^{\prime}}.$
We then have
$\frac{\mu}{6}-\widetilde{p}_{g}\geq\frac{a^{\prime}+b^{\prime}}{12}(k-1)\geq\frac{1}{6}.$
This concludes the proof. ∎
4.6.
We are now in a position to address the conjecture for the remaining quasi-homogeneous singularities in dimension one, which reduce to Proposition 4.4. For these cases, we don’t state the explicit expressions for the spectral genera, since they quickly become unmanageable.
Corollary 4.6.
The strong form of the conjecture is true for the singularities of the form (4.6) and (4.7).
Proof.
As in the proof of Proposition 4.4, we indicate the main steps of the argument. We begin with the case $x(x^{a}+y^{b})=0$. In this case, $\widetilde{p}_{g}$ is given by the sum
(4.10) $\sum\left(1-\frac{x}{a+1}-\frac{ay}{(a+1)b}\right),$
over the lattice points $(x,y)$ with $x>0$, $y>0$, in the interior of the triangle with vertices $(0,0)$, $(a+1,0)$, $(1,b)$. Actually, such points necessarily lie either in the interior of the triangle $T$ with vertices $(1,0)$, $(a+1,0)$ and $(1,b)$, or on the open edge joining $(1,0)$ and $(1,b)$, as in the following picture:
[…]
The triangle $T$, colored in pink, can be translated by one unit to the left so that, after the corresponding change of variables, the evaluation of (4.10) on the interior lattice points of $T$ reduces to the case treated in Proposition 4.4. The contribution of the lattice points on the open edge between $(1,0)$ and $(1,b)$, represented by the dashed line, reduces to a sum of consecutive integers. These computations are then combined with the value of the Milnor number, now given by $\mu=(a+1)(b-1)+1$. One concludes by inspection of the obtained expressions that $\mu/6-\widetilde{p}_{g}\geq 1/6$.
Next for the singularity $xy(x^{a}+y^{b})=0$. In this case, the spectral genus is given by the sum
(4.11) $\sum\left(1-\frac{bx}{(a+1)(b+1)-1}-\frac{ay}{(a+1)(b+1)-1}\right),$
over the lattice points $(x,y)$ either in the interior of the triangle $T$ with vertices $(1,1)$, $(a+1,1)$ and $(1,b+1)$, or on the edge joining $(1,1)$ and $(a+1,1)$, or on the edge joining $(1,1)$ and $(1,b+1)$. Notice that $(1,1)$ is one such point. This is represented in the following picture, where $T$ is again colored in pink:
[…]
The triangle can be translated so that $(1,1)$ is sent to the origin. The evaluation of (4.11) on the interior points of $T$ is then covered by Proposition 4.4. The sums on lattice points on the edges reduce to sums of consecutive integers. The result of the computation is then combined with the value of the Milnor number $\mu=(a+1)(b+1)$, and an examination of the expression yields again the bound $\mu/6-\widetilde{p}_{g}\geq 1/6$. ∎
5. Irreducible curve singularities
In this section, we consider the conjecture for an irreducible germ of a plane curve singularity, defined by $f\colon(\mathbb{C}^{2},0)\to(\mathbb{C},0)$. In this case, the spectrum can be described explicitly in terms of Puiseux pairs. This is discussed in an unpublished paper of M. Saito [30], whose presentation and notation we follow. The same result can be derived from the work of Schrauwen–Steenbrink–Stevens [31], as discussed by Némethi in [24].
5.1.
After possibly changing variables, the equation $f(x,y)=0$ is equivalent to a Puiseux series representation for $y$ in terms of $x$:
$\begin{split}y=&\sum_{1\leq i\leq[k_{1}/n_{1}]}c_{0,i}x^{i}+\sum_{0\leq i\leq[k_{2}/n_{2}]}c_{1,i}x^{(k_{1}+i)/n_{1}}\\ &+\sum_{0\leq i\leq[k_{3}/n_{3}]}c_{2,i}x^{k_{1}/n_{1}+(k_{2}+i)/n_{1}n_{2}}+\cdots+\sum_{i\geq 0}c_{g,i}x^{k_{1}/n_{1}+k_{2}/n_{1}n_{2}+\cdots+(k_{g}+i)/n_{1}\cdots n_{g}}.\end{split}$
The pairs $(k_{i},n_{i})$ are called Puiseux paris, and they fulfill the following properties:
  1. (1)
    $k_{i}$ and $n_{i}$ are coprime integers.
  2. (2)
    $n_{i}>1$.
  3. (3)
    $k_{1}>n_{1}$ and $k_{i}>k_{i-1}n_{i}$, for $i\geq 2$. In particular, $k_{i}>1$ for all $i$.
In addition, we introduce positive integers $w_{i}$, defined recursively by
$w_{1}=k_{1},\quad w_{i}=n_{i-1}n_{i}w_{i-1}+k_{i}\quad\text{for}\quad i\geq 2.$
Hence, for $i\geq 2$ we can write
(5.1) $w_{i}=\left(\sum_{j=1}^{i-1}k_{j}n_{j}(n_{j+1}\cdots n_{i-1})^{2}\right)n_{i}+k_{i}.$
For the sake of clarity, we stress that the term of index $j=i-1$ in the sum is understood to be $k_{i-1}n_{i-1}$. We also observe that, because the integers $k_{i}$ and $n_{i}$ are coprime, the same holds for $w_{i}$ and $n_{i}$.
5.2.
We next express the Milnor number and the spectral genus in terms of the previous quantities.
Lemma 5.1.
Let $f\colon(\mathbb{C}^{2},0)\to(\mathbb{C},0)$ define an irreducible germ of a plane curve singularity. Let the notation be as above, and define $n_{i}^{\prime}=n_{i+1}\cdots n_{g}$, with the convention $n_{g}^{\prime}=1$. Then:
  1. (1)
    The Milnor number is given by
    $\mu=\sum_{i=1}^{g}(n_{i}-1)(w_{i}-1)n_{i}^{\prime}.$
  2. (2)
    The spectral genus is given by
    $\widetilde{p}_{g}=\sum_{i=1}^{g}\sum\left(1-\frac{1}{n_{i}^{\prime}}\left(k+\frac{x}{n_{i}}+\frac{y}{w_{i}}\right)\right),$
    where the second sum runs over the integers $k\geq 0$, $x>0$ and $y>0$, satisfying $k<n_{i}^{\prime}$ and $x/n_{i}+y/w_{i}<1$.
Proof.
See Némethi [24, Section 3], and in particular Theorem 3.1 therein, and M. Saito [30, Theorem 1.5 & Section 5]. For the spectral genus, we refer to §2.3 above for the expression in terms of Saito’s convention. ∎
Corollary 5.2.
With the assumptions and notation as above, we have
$\frac{\mu}{6}-\widetilde{p}_{g}=\frac{1}{12}\sum_{i=1}^{g}S_{i},$
where $S_{i}=S_{i}^{+}-S_{i}^{-}$ and
$S_{i}^{+}=\frac{(n_{i}-1)(w_{i}-1)(n_{i}+w_{i}+1)}{n_{i}w_{i}},\quad S_{i}^{-}=(n_{i}-1)(w_{i}-1)(n_{i}^{\prime}-1).$
Proof.
This is an elementary computation using the expressions provided in Lemma 5.1 and applying Proposition 4.4 in the case when $a$ and $b$ are coprime. ∎
5.3.
We are now in a position to state and prove the main theorem of this section.
Theorem 5.3.
The strong form of the conjecture holds for germs of irreducible plane curve singularities. More precisely,
$\frac{\mu}{6}-\widetilde{p}_{g}\geq\frac{1}{12}S_{1}^{+}\geq\frac{1}{6},$
and the first inequality is strict if there are two or more Puiseux pairs.
Proof.
Observe that $S_{g}^{-}=0$, since $n_{g}^{\prime}=1$. Hence, the case $g=1$ is trivial and we may assume that $g\geq 2$. We have to prove
$\sum_{i=2}^{g}S_{i}^{+}-\sum_{i=1}^{g-1}S_{i}^{-}>0.$
First of all, for $i\geq 2$, we provide a lower bound for $S_{i}^{+}$ which is simpler to deal with. For this, using that $k_{i}>1$ in (5.1), and that $(n_{i}+w_{i}+1)/w_{i}>1$, we see that
(5.2) $S_{i}^{+}>T_{i}^{+}:=(n_{i}-1)\sum_{j=1}^{i-1}k_{j}n_{j}(n_{j+1}\cdots n_{i-1})^{2},$
where the summand of index $j=i-1$ is understood to be $k_{i-1}n_{i-1}$.
Next, for all $i$, $S_{i}^{-}$, has the following simple upper bound:
(5.3) $\begin{split}S_{i}^{-}\leq T_{i}^{-}:&=(n_{i}-1)(n_{i}^{\prime}-1)w_{i}\\ &=(n_{i}-1)(n_{i+1}\cdots n_{g}-1)\left(k_{i}+\sum_{j=1}^{i-1}k_{j}n_{j}(n_{j+1}\cdots n_{i-1})^{2}n_{i}\right).\end{split}$
Using the above bounds for $S_{i}^{+}$ and $S_{i}^{-}$, we can write
(5.4) $\sum_{i=2}^{g}S_{i}^{+}-\sum_{i=1}^{g-1}S_{i}^{-}>\sum_{i=2}^{g}T_{i}^{+}-\sum_{i=1}^{g-1}T_{i}^{-}=a_{1}k_{1}+\cdots+a_{g}k_{g},$
and it is enough to prove that the coefficients $a_{i}$ are positive. We will discuss the coefficient $a_{1}$. The other coefficients are dealt with similarly.
We denote by $P_{i}$ the coefficient of $k_{1}$ in $T_{i}^{+}$. Similarly, we denote by $N_{i}$ the coefficient of $k_{1}$ in $T_{i}^{-}$. Inspecting the expanded expressions (5.2) and (5.3) for $T_{i}^{+}$ and $T_{i}^{-}$, we find the following. For the coefficients $P_{i}$, we have
$P_{2}=n_{1}(n_{2}-1)$
and
$P_{i}=n_{1}(n_{2}\cdots n_{i-1})^{2}(n_{i}-1)$
for $i\geq 3$. For the coefficients $N_{i}$, we have
$N_{1}=(n_{1}-1)(n_{2}\cdots n_{g}-1)$
and
$N_{2}=n_{1}n_{2}(n_{2}-1)(n_{3}\cdots n_{g}-1),$
while for $i\geq 3$ we decompose $N_{i}=N_{i,a}+N_{i,b}$, where
$N_{i,a}=n_{1}(n_{2}\cdots n_{i})^{2}n_{i+1}\cdots n_{g}-n_{1}(n_{2}\cdots n_{i-1})^{2}n_{i}\cdots n_{g}$
and
$N_{i,b}=n_{1}(n_{2}\cdots n_{i-1})^{2}n_{i}-n_{1}(n_{2}\cdots n_{i})^{2}.$
The sum of the coefficients $N_{i,a}$ is a telescopic sum, with value
$\sum_{i=3}^{g-1}N_{i,a}=n_{1}(n_{2}\cdots n_{g-1})^{2}n_{g}-n_{1}n_{2}^{2}n_{3}\cdots n_{g}.$
Adding the contributions of $P_{2}$, $N_{1}$ and $N_{2}$, we obtain
(5.5) $P_{2}-N_{1}-N_{2}-\sum_{i=3}^{g-1}N_{i,a}=n_{2}\cdots n_{g}+n_{1}n_{2}^{2}-n_{1}(n_{2}\cdots n_{g-1})^{2}n_{g}-1.$
Next, we consider the coefficients $P_{i}$ together with the coefficients $N_{i,b}$, for $13\leq i\leq g-$ 1. More precisely,
$P_{i}-N_{i,b}=n_{1}(n_{2}\cdots n_{i})^{2}-n_{1}(n_{2}\cdots n_{i-1})^{2},$
which again gives rise to a telescopic sum:
$\sum_{i=3}^{g-1}(P_{i}-N_{i,b})=n_{1}(n_{2}\cdots n_{g-1})^{2}-n_{1}n_{2}^{2}.$
Adding the coefficient $P_{g}=n_{1}(n_{2}\cdots n_{g-1})^{2}(n_{g}-1)$, we find
(5.6) $\sum_{i=3}^{g-1}(P_{i}-N_{i,b})+P_{g}=n_{1}(n_{2}\cdots n_{g-1})^{2}n_{g}-n_{1}n_{2}^{2}.$
To conclude, we add (5.5) and (5.6), which yields the coefficient $a_{1}$ in (5.4):
$a_{1}=n_{2}\cdots n_{g}-1>0.$
This completes the proof. ∎
6. Generic singularities with large Newton polyhedra
In this section, we consider isolated hypersurface singularities which are non-degenerate and convenient in the sense of Kouchnirenko [20]. [^2] In op. cit., which is written in French, convenient is called commode. These conditions are formulated in terms of the Newton polyhedron. The non-degeneracy condition is shown to be generic for the Zariski topology. Therefore, we will refer to such singularities as generic. Convenient means that the Newton polyhedron intersects all the coordinate hyperplanes. We refer to op. cit. for details.
6.1.
Consider now a convenient and generic isolated hypersurface singularity defined by $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$. The formula of Kouchnirenko [20, Théorème I] states that
(6.1) $\mu=(n+1)!\operatorname{vol}(\Gamma_{-})-n!\operatorname{vol}_{n}(\Gamma_{-})+\ldots+(-1)^{n}\operatorname{vol}_{1}(\Gamma_{-})+(-1)^{n+1}.$
Here, $\operatorname{vol}$ is the standard volume in $\mathbb{R}^{n+1}$, and $\operatorname{vol}_{k}(\Gamma_{-})$ refers to the volume of the lower Newton polyhedron intersected with all the coordinate subspaces of dimension $k$, with respect to the standard Lebesgue measure. In particular, $\operatorname{vol}_{n}(\Gamma_{-})$ is the volume of $\Gamma_{-}$ intersected with the coordinate hyperplanes.
The spectral genus can also be described explicitly in terms of the Newton polyhedron. Let $\Gamma_{c}$ be the compact Newton boundary. It is defined by a homogeneous, concave, piece-wise linear function $\phi$, which takes the value 1 on the faces of $\Gamma_{c}$. The lower Newton polyhedron can then be presented as $\Gamma_{-}=\{x\in\mathbb{R}_{+}^{n+1}\mid\phi(x)\leq 1\}$. Then, by [29], in particular using the formulas on the first page, it follows that the spectral genus is given by
(6.2) $\widetilde{p}_{g}=\sum_{x\in\Gamma_{-}^{\circ}\cap\mathbb{Z}^{n+1}}(1-\phi(x)),$
where the sum runs over the interior lattice points of $\Gamma_{-}$. Here we have taken into account the different normalization as recalled in (2.4). This sum is analogous to (4.2) but appears here in an a priori different context.
6.2.
In order to estimate an asymptotic form of the sum (6.2), we will apply a rough version of the Euler–Maclaurin formula for polyhedra. For the statement, we adopt the following convention for Lebesgue measures. Let $P$ be a convex lattice polyhedron in $\mathbb{R}^{d}$, and $Q$ a face of it. Let $\langle Q\rangle_{\mathbb{R}}\subseteq\mathbb{R}^{d}$ be the real vector space parallel to $Q$. Then we denote by $dx$ the Lebesgue measure on $\langle Q\rangle_{\mathbb{R}}$ which gives volume one to the fundamental mesh of the lattice $\mathbb{Z}^{d}\cap\langle Q\rangle_{\mathbb{R}}$.
Lemma 6.1 (Berline–Vergne).
Let $P\subset\mathbb{R}^{d}$ be a compact convex lattice polyhedron, and let $h$ be a smooth function on $P$. Then, there exists an asymptotic expansion
[…]
Proof.
This is an application of the local Euler–Maclaurin formula of Berline–Vergne [2, Section 5.4, Theorem 5]. The dominant term of the expansion is already given in loc. cit. They discuss the subdominant term in Section 5.5. Notice also that in [2] one supposes that $h$ is a smooth function on $\mathbb{R}^{d}$ with compact support. Since the sum and the first terms of the expansion depend only on the values of $h$ in $P$, which is compact, we can smoothly extend $h$ outside of $P$, still with compact support, for which the formula applies. ∎
6.3.
In general, the integrals appearing in the previous lemma can be evaluated by applying results of Brion [7, Section 3.2], see also [8, Corollary 6.1.10]. In our setting, it will be enough to have the following.
Lemma 6.2.
Let $P$ be a $d$-dimensional simplex in $\mathbb{R}^{d}$, with vertices $0,u_{1},\ldots,u_{d}$. Let $\lambda\in\mathbb{R}^{d}$ be such that $\langle u_{i},\lambda\rangle=1$ for all $i$. Then
[…]
Proof.
The formula is equivalent to the statement that
[…]
which concludes the computation. ∎
Theorem 6.3.
Let $f(x_{0},\ldots,x_{n})=0$ be a convenient and generic isolated singularity. Then for $k$ sufficiently large, the isolated singularity $f(x_{0}^{k},\ldots,x_{n}^{k})=0$ satisfies the strong form of the conjecture.
Proof.
First of all, notice that the singularity $f(x_{0}^{k},\ldots,x_{n}^{k})=0$ is still generic and convenient. Hence, the discussion §6.1 applies to it. If we denote the corresponding Milnor number by $\mu(k)$, then by (6.1) we have the asymptotic expansion
(6.4) $\frac{\mu(k)}{(n+2)!}=\frac{k^{n+1}}{n+2}\operatorname{vol}(\Gamma_{-})-\frac{k^{n}}{(n+1)(n+2)}\operatorname{vol}_{n}(\Gamma_{-})+O(k^{n-1}).$
Similarly, we consider
(6.5) $\widetilde{p}_{g}(k)=\sum_{x\in(k\Gamma_{-}^{\circ})\cap\mathbb{Z}^{n+1}}\left(1-\phi(x/k)\right),$
where $\phi$ is the piece-wise linear function determining the lower Newton polyhedron $\Gamma_{-}$. To estimate (6.5), we apply Lemma 6.1. For this, we first decompose $\Gamma_{-}$ into simplices with 0 as a vertex and the rest of vertices in $\mathbb{Z}^{n+1}\cap\Gamma_{c}$. These define a simplicial complex, whose top dimensional simplices we denote by $\sigma_{i}$, and whose codimension 1 simplices we denote by $\tau_{j}$. On each simplex, the function $\phi$ is linear. A first application of the lemma gives
$\widetilde{p}_{g}(k)=\sum_{i}\sum_{x\in k\sigma_{i}\cap\mathbb{Z}^{n+1}}(1-\phi(x/k))-\sum_{j}\sum_{x\in k\tau_{j}\cap\mathbb{Z}^{n+1}}(1-\phi(x/k))+O(k^{n-1}).$
Applying Lemma 6.1 once again, together with Lemma 6.2, one straightforwardly deduces that
(6.6) $\widetilde{p}_{g}(k)=\frac{k^{n+1}}{n+2}\operatorname{vol}(\Gamma_{-})-\frac{k^{n}}{2(n+1)}\operatorname{vol}_{n}(\Gamma_{-})+O(k^{n-1}).$
Combining (6.4) and (6.6) we find that
$\frac{\mu(k)}{(n+2)!}-\widetilde{p}_{g}(k)=k^{n}\frac{n}{2(n+1)(n+2)}\operatorname{vol}_{n}(\Gamma_{-})+O(k^{n-1}).$
For big enough $k$, the leading term, which is positive, dominates and this concludes the proof of the theorem. ∎
Corollary 6.4.
With the assumptions of Theorem 6.3, we have
$\frac{\widetilde{p}_{g}(k)}{\mu(k)}\nearrow\frac{1}{(n+2)!}\quad\text{ as }\quad k\to+\infty,$
where $\mu(k)$ and $\widetilde{p}_{g}(k)$ are the Milnor number and spectral genus of $f(x_{0}^{k},\ldots,x_{n}^{k})=0$.
7. Relation to other conjectures
In this section, we discuss the relationship between our conjecture and other conjectural statements, due to K. Saito and C. Hertling, which focus on the distribution of the spectral numbers.
7.1. K. Saito’s conjecture
In [27], K. Saito proposed a conjectural density describing the distribution of the spectral numbers of families of isolated hypersurfaces singularities, as the singularity gets worse. In such situations, K. Saito’s conjecture entails in particular that the inequality in our conjecture is asymptotically an equality, and hence our conjecture is sharp.
7.2.
A family $\mathcal{F}$ of $n$-dimensional isolated hypersurface singularities, defined by germs of holomorphic functions $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$, will be called degenerating if it is endowed with a filter and the function $\mu\colon\mathcal{F}\to\mathbb{R}$ induced by the Milnor number tends to infinity. That is,
$\lim_{\mathcal{F}}\mu=+\infty.$
The filter is a technical device which allows us to talk about limits in a rigorous fashion. We informally see $\mathcal{F}$ as a family of singularities whose Milnor numbers converge to infinity.
7.3.
To state K. Saito’s conjecture, let $f$ be a germ of a holomorphic function defining an isolated hypersurface singularity at the origin. Recall the notion of the associated spectral numbers, that we now take in the form $10<\alpha_{1}^{\prime}\leq\ldots\leq\alpha_{\mu}^{\prime}<n+$ 1. We define the corresponding spectral probability measure as
(7.1) $\delta_{f}=\frac{1}{\mu}\sum_{j}\delta_{\alpha_{j}^{\prime}}.$
For a fixed $n$, consider the measure $N(s)ds$ on $[0,n+1]$ for which
[…]
where the $x_{i}\in[0,1]$. Notice this is a push-forward measure. Indeed, denote by $\nu$ the measure on $[0,1]^{n+1}$ induced by the Lebesgue measure on $\mathbb{R}^{n+1}$, and introduce the addition function $\Sigma(x_{0},\ldots,x_{n})=\sum x_{i}:[0,1]^{n+1}\to[0,n+1]$. Then $N(s)ds=\Sigma_{\ast}(\nu)$.
Conjecture (K. Saito).
For suitable degenerating families of singularities $\mathcal{F}$, we have the convergence
(7.2) $\lim_{\mathcal{F}}\delta_{f}(s)=N(s)ds,$
of probability measures on $\mathbb{R}$, in the strong sense.
It is part of the conjecture to find meaningful families of singularities for which the statement holds. We recall that for probability measures on $\mathbb{R}$, strong convergence means convergence of the measures of any measurable set. This is equivalent to the convergence of the integrals of any bounded measurable function. The convergence in the strong sense is important in the applications below, where we need to integrate functions which are not continuous, but at least measurable.
7.4.
The Fourier transform of the Dirac distribution $\delta_{f}$ can be expressed in terms of the spectral polynomial of $f$ (2.5). By considering cases when the latter is known, one can prove K. Saito’s conjecture in the following situations:
  1. (1)
    Quasi-homogeneous singularities, with weights $w_{0},\ldots,w_{n}$ converging to 0, proven by Saito [27, §3.7, Example 1].
  2. (2)
    Irreducible plane curve singularities, with Puiseux pairs $(k_{1},n_{1}),\ldots,(k_{g},n_{g})$ and $n_{g}\to+\infty$, proven by Saito [27, §3.9, Example 3].
  3. (3)
    For convenient, generic singularities, with large Newton polyhedra as in Section 6 above, proven by Almirón–Schulze [1, Theorem 1.1].
7.5.
By applying (7.2) to well-chosen functions, one can derive necessary conditions on a family $\mathcal{F}$ to satisfy K. Saito’s conjecture. We state some of these special features.
Proposition 7.1.
Let $\mathcal{F}$ be a degenerating family for which K. Saito’s conjecture holds. Then:
  1. (1)
    The Hodge numbers of the Milnor fibers satisfy
    $\lim_{\mathcal{F}}\dim\operatorname{Gr}_{F}^{p}H^{n}(\operatorname{Mil}_{f})=+\infty,\quad\text{for}\quad 0\leq p\leq n.$
  2. (2)
    The minimal spectral value $\alpha_{1}^{\prime}$ satisfies
    $\lim_{\mathcal{F}}\alpha_{1}^{\prime}=0.$
Proof.
For the first point, we have
[…]
where $A_{p}$ is the region in $[0,1]^{n+1}$ defined by $n-p\leq\sum x_{i}\leq n-p+1$. Because the integral over $A_{p}$ is strictly positive and $\mu$ converges to +$\infty$+ ∞ along $\mathcal{F}$, the same must hold for $\dim\operatorname{Gr}_{F}^{p}H^{n}(\operatorname{Mil}_{f})$.
For the second point, fix $\varepsilon>0$. We have
[…]
where we used that $A_{n}$ is the standard simplex in +$\mathbb{R}^{n+1}_{+}$, whose volume is $1/(n+1)!$. That is, for every $\varepsilon^{\prime}>0$, we can find an element of the filter $\mathcal{F}_{\varepsilon^{\prime}}$ such that, for all $f\in\mathcal{F}_{\varepsilon^{\prime}}$,
|#$\left|\frac{\#\{\alpha_{j}^{\prime}<\varepsilon\}}{\mu}-\frac{\varepsilon^{n+1}}{(n+1)!}\right|<\varepsilon^{\prime}.$
In particular, for $\varepsilon^{\prime}=\varepsilon^{n+1}/2(n+1)!$, we infer that the set #$\#\{\alpha_{j}^{\prime}<\varepsilon\}$ is non-empty for all $f\in\mathcal{F}_{\varepsilon^{\prime}}$. Consequently, $\alpha_{1}^{\prime}<\varepsilon$. Because $\alpha_{1}^{\prime}>0$, this concludes the proof. ∎
7.6.
For the geometric genus and the spectral genus, K. Saito’s conjecture leads to the following expectations.
Proposition 7.2.
Assume that K. Saito’s conjecture holds for some degenerating family of singularities $\mathcal{F}$. Then:
  1. (1)
    The geometric genus satisfies
    $\lim_{\mathcal{F}}\frac{p_{g}}{\mu}=\frac{1}{(n+1)!}.$
  2. (2)
    The spectral genus satisfies
    (7.4) $\lim_{\mathcal{F}}\frac{\widetilde{p}_{g}}{\mu}=\frac{1}{(n+2)!}.$
  3. (3)
    In particular, we have
    $\lim_{\mathcal{F}}\frac{\widetilde{p}_{g}}{p_{g}}=\frac{1}{n+2}.$
Proof.
The first item follows from (7.3) in the case $n=p$, together with the evaluation of the volume of the standard simplex $A_{n}$.
For the second item, we first observe that
$\frac{\widetilde{p}_{g}}{\mu}=\int_{0}^{1}(1-s)\delta_{f}(s),$
according to (2.4). Under K. Saito’s conjecture, this converges to
[…]
Since the volume of $A_{n}$ is $1/(n+1)!$, by Lemma 6.2 we conclude that the value of the integral is $1/(n+2)!$.
The third point of the proposition is a combination of the first and the second points. ∎
The second item of the proposition should be compared to Proposition 4.2 and Corollary 6.4, to the effect that (7.4) not only holds in the corresponding cases, but the limit is approached from below. Also, we see that Proposition 4.2 and Corollary 6.4 are compatible with the known cases of K. Saito’s conjecture reviewed in §7.4 above.
7.7. Hertling’s conjecture
While K. Saito’s conjecture describes the distribution of the spectral numbers of an isolated singularity, Hertling’s conjecture focuses on the variance. We discuss the relationship to our conjecture. Combined with K. Saito’s conjecture, we conclude that Hertling’s conjecture does not seem to imply our conjecture.
7.8.
Consider the spectral numbers of an isolated hypersurface singularity, given in the form $-1<\alpha_{1}\leq\ldots\leq\alpha_{\mu}\leq n$. Recall these are symmetric with respect to $\alpha\mapsto n-1-\alpha$. That is, we have the relationship $\alpha_{j}+\alpha_{\mu-j+1}=n-1$. Hence, the mean value of the spectral values is $(n-1)/2$. In [16], Hertling proposed a bound for the variance.
Conjecture (Hertling).
The spectral numbers of an isolated hypersurface singularity, taken in the interval $(-1,n)$, satisfy
(7.5) $\frac{1}{\mu}\sum_{j}\left(\alpha_{j}-\frac{n-1}{2}\right)^{2}\leq\frac{\alpha_{\mu}-\alpha_{1}}{12}.$
Hertling’s conjecture is known for quasi-homogeneous singularities. In this case, the predicted inequality is in fact an equality. This was proven by Hertling himself in [16]. An elementary proof was later found by Dimca [10]. The case of curve singularities has been addressed by several authors. In the unpublished article [30], M. Saito proved the case of irreducible plane curve singularities. In [4], Brélivet settled the case of plane curves with convenient, non-degenerate Newton polygon. Later, in the unpublished article [5], he considered general plane curve singularities.
Theorem 7.3 (Brélivet, M. Saito).
Hertling’s conjecture holds for isolated plane curve singularities.
The previous statement has the following consequence for our conjecture.
Corollary 7.4.
For plane curve singularities, the weak form of the conjecture is true if $\alpha_{\mu}<2/3$.
Proof.
For plane curve singularities, the spectral numbers normalized in $(-1,1)$ are invariant under the symmetry $x\mapsto-x$. In particular, $\alpha_{\mu}=-\alpha_{1}$. Hence, combining (7.5) with the Cauchy–Schwarz inequality, we derive
$\widetilde{p}_{g}=\sum_{0<\alpha_{j}<1}\alpha_{j}\leq\mu\sqrt{\frac{\alpha_{\mu}}{24}}.$
The claim is a straightforward consequence of the latter inequality. ∎
7.9.
Suppose now that we are in the setting of a family $\mathcal{F}$ of plane curve singularities for which K. Saito’s conjecture holds. We can then apply Proposition 7.1 (2). Recalling that $\alpha_{1}^{\prime}=\alpha_{1}+1$ and $\alpha_{1}=-\alpha_{\mu}$, we find
$\lim_{\mathcal{F}}\alpha_{\mu}=1.$
Therefore, we see that the criterion provided by the corollary, which requires $\alpha_{\mu}<2/3$, will apply at most for some exceptional singularities of the family $\mathcal{F}$, but not in general.
Funding
The first author is supported by the Swedish Research Council, VR grant 2021-03838 "Mirror symmetry in genus one". The second author is supported by the Knut och Alice Wallenberg foundation "Guest researcher program".
Acknowledgements
We thank Omid Amini, Jean-Benoît Bost, José Ignacio Burgos Gil, Lars Halvard Halle, Thomas Mordant, Claude Sabbah, Kyoji Saito and Jan Stevens for discussions relating to this article. We are in particular grateful to Ken-Ichi Yoshikawa for many discussions and the hospitality in Kyoto, and for many insights into the analytic torsion. The possibility of the continuity of the analytic torsion on Hilbert schemes was suggested by him. The second author wishes to thank the Department of Mathematics at Chalmers University of Technology and the University of Gothenburg for the hospitality too.